# Rate, Explain and Cite (REC): Enhanced Explanation and Attribution in Automatic Evaluation by Large Language Models

### v1

@@ Acknowledgments
Appendix A Various RAG Citation Modes that the REC Models Support
Here are example output templates of the different citation modes that our model supports and has been evaluated on. Users can easily process information from the JSON objects into the corresponding natural language format they desire. For all of these output styles, here’s the example response and context that we’ve used to feed into the model.
@@ Appendix A Various RAG Citation Modes that the REC Models Support
A.1 Post-fix citation
@@
A.2 Post-fix citation with snippet extracted verbatim from retrieved context
@@
A.3 Inline citation with claims extracted verbatim from LLM generated response
@@
A.4 Inline citation with snippet
@@
A.5 Prompt for Inline Citation
Below is the prompt that we used to generate inline citations. This can be modified to generate responses in each of the different RAG citation modes shown in the sections above.
Appendix B REC-Data Details
For general instruction tasks, we include summarization on ABCD (Chen et al., 2021), a dataset containing customer support dialogue transcripts, in five languages. For pointwise evaluation, we leverage HelpSteer2 (Wang et al., 2024c) and perform synthetic data generation (Section 3.2) on the (dialogue, summary) pairs from ABCD, and QA pairs from FloDial. For pairwise and open-ended evaluation tasks, we use Auto-j (Li et al., 2023a), Shepherd (Wang et al., 2023b), Skywork  [^4] https://huggingface.co/datasets/Skywork/Skywork-Reward-Preference-80K-v0.1, HelpSteer2555https://huggingface.co/datasets/nvidia/HelpSteer2, OffsetBias [^6] https://huggingface.co/datasets/NCSOFT/offsetbias, and Code Preference [^7] https://huggingface.co/datasets/Vezora/Code-Preference-Pairs. For citations, we perform synthetic data generation (Section 3.2) on FloDial for both RAG citations and content quality citations, and on ABCD, Auto-j, Shepherd, and Email Thread Summary  [^8] https://www.kaggle.com/datasets/marawanxmamdouh/email-thread-summary-dataset/data for content quality citations.
@@ Appendix B REC-Data Details
Appendix C Synthetic Data Generation Prompt Example
C.1 Pointwise Evaluation ("Faithfulness metric")
@@
C.2 Rating, Explanation, and Citation Evaluation Prompt
Appendix D Instruction for Claim + Citations Labeling
We’ve asked an LLM to evaluate another LLM’s ’Answer’ for a summarization task given a ’Query’ and ’Context’. The evaluator LLM has generated explanations for its evaluation with supporting citations, as shown in the ’metric_full_explain’, ’Claim’ (parts extracted from the explanation) and ’Citations’ columns in a file separately.
D.1 Instructions
You have two tasks to complete:
  1. 1.
    Evaluate the quality of the ’Claim’ generated by the evaluator LLM based on its corresponding ’Metric’. For each claim, first identify which metric it’s trying to explain against by looking at the ’Metric’ column, and refer to the corresponding label guideline for that metric below. Fill in your answer in ’Claim_correct’. Note, you don’t need to read the ’metric_full_explain’ column to decide if the claim is correct or not.
    • Faithfulness:
      • Choose ’Yes’: A claim for faithfulness is correct if it follows ALL the three criteria: (1) it is a correct judgement, (2) it correctly justifies that the answer either contains only accurate information mentioned in the context or points out the invented facts, and (3) it doesn’t contain irrelevant or extra information in addition to the justification itself.
      • Choose ’No’: otherwise.
    • Instruction-following:
      • Choose ’Yes’: A claim for instruction-following is correct if it follows ALL the three criteria: (1) it is a correct judgement, (2) it correctly justifies why (or why not) the answer is relevant to the query and follows what’s asked in the query, not performing any extra tasks or repeating the same task more than once, and (3) it doesn’t contain irrelevant or extra information in addition to the justification itself.
      • Choose ’No’: otherwise.
    • Coherence:
      • Choose ’Yes’: A claim for coherence is correct if it follows ALL the three criteria: (1) it is a correct judgement, (2) it correctly justifies why (or why not) the answer is easy to understand and free of grammar errors, and (3) it doesn’t contain irrelevant or extra information in addition to the justification itself.
      • Choose ’No’: otherwise.
    • Completeness:
      • Choose ’Yes’: A claim for completeness is correct if it follows ALL the three criteria: (1) it is a correct judgement, (2) it correctly justifies that the answer either covers the desired content to understand the result or points out the missing information that can make the result confusing or misleading, and (3) it doesn’t contain irrelevant or extra information in addition to the justification itself.
      • Choose ’No’: otherwise.
  2. 2.
    [Only proceed to this step if ’Claim_correct’ = ’Yes’] Evaluate the quality of the ’Citations’ generated by the evaluator LLM, and correct the citations if needed.
    For each claim, there’s a list of citations consisting of verbatim sentences extracted from ’Query’, ’Context’, and ’Answer’, with source specified in the parentheses, as evidence to support the claim. Please evaluate the correctness of the generated citations as below:
    • Choose ’Yes’ for ’Citations_correct’: The citations are correct if it follows ALL the three criteria: (1) the sources in the parentheses are correct, (2) no sentences are rewritten/paraphrased, and (3) they are all relevant evidence to support the claim, and there are no missing relevant sentences not cited from either ’Query’, ’Context’, or ’Answer’. You don’t have to do anything for ’Revised_citations’ in this case.
    • Choose ’No’ for ’Citations_correct’: If any of the three criteria is not met, choose ’No’ for ’Citations_correct’, and revise the generated citations to provide a correct list of citations (following the generated format) in ’Revised_citations’.
Appendix E Prompt for LLM-AggreFact Eval

### v2

@@ Acknowledgments
Appendix A Content Quality Citation Modes and the Evaluation Prompt
Here we provide the evaluation prompt template used in Figure 1, the raw autoevaluator output, and a final output example in an alternative citation mode for content quality evaluation.
A.1 Content Quality Evaluation Prompt
A.2 Content Quality Citation Modes
Using the same input example as in Figure 1, we show the raw autoevaluator output that allows users to further process according to either of the two citation modes, “post-fix with context snippet” and “inline with context snippet”, to fit their needs. With the final output in “inline with context snippet” mode already presented in Figure 1, here we show another possible final output format in “post-fix with context snippet” mode.
A.2.1 Raw Autoevaluator Output
Citations are extracted verbatim from the input context to support statements in the autoevaluator’s explanation.
"statements": [ {   "statement_string": "The answer covers all   the key points of the conversation,   including the customer’s question, the   agent’s response, and the resolution.",   "citations": ["Summarizing key points   of the conversation including customer   issue and resolution.", "Customer: Is it   possible to have my items gift wrapped?",   "Agent: We offer gift wrapping but it   cost \$4.99 per item"] }, {   "statement_string": "It also mentions the   customer’s reason for asking about gift   wrapping and the agent’s polite closing.",   "citations": ["Customer: Oh great. My   Bestie’s birthday is coming up, that’s   the reason for my questions. Thanks",   "Agent: All right. thanks for   reaching out and have a great day!"]  }]
A.2.2 Final Output Example in “post-fix with context snippet” Citation Mode
The answer covers all the key points of theconversation, including the customer’s question,the agent’s response, and the resolution. It alsomentions the customer’s reason for asking aboutgift wrapping and the agent’s polite closing.[1][2][3][4][5][1]: “Summarize key points of the conversationincluding customer issue and resolution.“[2]: “Customer: Is it possible to have my itemsgift wrapped?”[3]: “Agent: We offer gift wrapping but it costs\$4.99 per item.”[4]: “Customer: Oh great. My Bestie’s birthdayis coming up, that’s the reason for my questions.Thanks.“[5]: “Agent: All right. thanks for reaching outand have a great day!” “
Appendix B Details of using REC Models for General RAG citations
Figure 2: Illustration of using REC models for general RAG citations.
We demonstrate how REC models can also be used for general RAG citations in Figure 2. In the example shown, we present this question “What is photosynthesis?” to another LLM, and collect retrieved articles with ids as the context, and the generated answer from another LLM to serve as inputs. The goal is for our LLM autoevaluator to provide citations extracted verbatim from the context to support the generation from another LLM. REC models support four different citation modes, and the details can also be found in this section.
@@ Appendix B Details of using REC Models for General RAG citations
B.1 RAG Citation Prompt
As an example, below is the prompt that we used to generate inline citations. This can be modified to generate responses in each of the different RAG citation modes shown in the following sections.
B.2 Post-fix Citation
@@
B.3 Post-fix Citation with Context Snippet
@@
B.4 Inline Citation
@@
B.5 Inline Citation with Context Snippet
@@
Appendix C REC-Data Details
For general instruction tasks, we include summarization on ABCD (Chen et al., 2021), a dataset containing customer support dialogue transcripts, in five languages. For pointwise evaluation, we leverage HelpSteer2 (Wang et al., 2024c) and perform synthetic data generation (Section 4.2) on the (dialogue, summary) pairs from ABCD, and QA pairs from FloDial. For pairwise and open-ended evaluation tasks, we use Auto-j (Li et al., 2023a), Shepherd (Wang et al., 2023b), Skywork  [^10] https://huggingface.co/datasets/Skywork/Skywork-Reward-Preference-80K-v0.1, HelpSteer2111111https://huggingface.co/datasets/nvidia/HelpSteer2, OffsetBias [^12] https://huggingface.co/datasets/NCSOFT/offsetbias, and Code Preference [^13] https://huggingface.co/datasets/Vezora/Code-Preference-Pairs. For citations, we perform synthetic data generation (Section 4.2) on FloDial for both RAG citations and content quality citations, and on ABCD, Auto-j, Shepherd, and Email Thread Summary  [^14] https://www.kaggle.com/datasets/marawanxmamdouh/email-thread-summary-dataset/data for content quality citations.
@@ Appendix C REC-Data Details
Appendix D Synthetic Data Generation Prompt Example
The prompts used to generate synthetic content quality and RAG citation data are the same as the ones shown in Appendices A and B.
D.1 Pointwise Evaluation ("Faithfulness metric")
@@
Appendix E Remark on Why The Three Metrics in ALCE are All Important
Regarding why the three criteria are equally important, we encourage the reader to consider the user’s perspective. An LLM’s citation response cannot be deemed good if it is lacking in any of the three criteria: fluency, answer correctness, and citation accuracy. To illustrate this, we aim to provide an example of a good citation response that scores well on all three criteria, as well as three examples that each fail in one of the criteria (assuming the same RAG QA setting as in Appendix B).
  • Example 1 (scoring high in all criteria): Photosynthesis is the process by which green plants and some other organisms utilize sunlight to synthesize their food. This remarkable process involves the conversion of carbon dioxide and water into glucose and oxygen, facilitated by chlorophyll [1233].
  • Example 2 (lacks fluency): Photosynthesis, it’s the way that green plants and also some other organisms they use (*not fluent) sunlight for making their food. This process, it’s interesting, converts carbon dioxide with water into glucose and oxygen, and chlorophyll helps with it [1233].
  • Example 3 (lacks correctness): Photosynthesis, it’s when green plants and other organisms use water (*wrong) to directly create sunlight (*wrong) as food. This process, it’s interesting, turns oxygen into glucose and carbon dioxide, with chlorophyll doing nothing important.
  • Example 4 (lacks citation accuracy): Photosynthesis is the process by which green plants and some other organisms utilize sunlight to synthesize their food. This remarkable process involves the conversion of carbon dioxide and water into glucose and oxygen, facilitated by chlorophyll [1422] (*wrong citation).
It is highly likely that a user would prefer the first example, which is well-balanced across all three criteria, as a good response. We hope this illustration better conveys our point.
Appendix F ABCD Content Quality Citation Human Evaluation Instruction
Annotation instruction: You will be evaluating attributed judgments made by an LLM autoevaluator for answers generated by another LLM against three metrics: instruction-following, completeness, and faithfulness. Your task is to evaluate LLM autoevaluator’s answer by completing the following THREE tasks: *you can skip if the output shows “invalid json”.
  • Metric answer correctness: Please label “yes”/“no” by evaluating whether the metric “answer” in autoevaluator’s output is correct or not, based on the provided conversation, task prompt, and summary.
  • Metric explanation correctness: Please label “yes”/“no’ by evaluating whether the “feedback” in autoevaluator’s output is correct or not, based on the provided conversation, task prompt, and summary. Only proceed to citation annotation (next task) if you label “yes” in this task, namely the explanation is correct.
  • Citation annotation: In “statements”, you will find different “statement_string” extracted from the “feedback”. Please provide citation for each “statement_string” by directly copying relevant sentences (Please cite full sentences, not substrings) from “task prompt” and “chat” columns (Please DON’T cite from “summary”). If the “statement_string” is hallucinated and doesn’t make sense in the first place, please put “[<halu>]” as the citation to signal.
Appendix G Prompt for LLM-AggreFact Eval